OpenStudy (anonymous):

Can any one help me with this question (A) Solve 3x^2-7x-13=0 (B) find the x-intercepts of f(x)= 3x^2-7x-13 (A) What are the solutions? x= (B) What are the x-intercepts?

OpenStudy (anonymous):

use quadratic equation here x=[-b-+sqrt(b^2 -4ac)]/2a x=[-b-+sqrt((-7)^2 -4(3)(-13))]/2(3) x1=(7+sqrt205)/6 ans... x2=(7-sqrt205)/6 ans...